OpenStudy (anonymous):

2/(3w)=2/15+12/(5w) show work

OpenStudy (gorv):

\[\frac{ 2 }{ 3*w }=\frac{ 2 }{ 15 }+\frac{ 12 }{ 5w }\]

OpenStudy (gorv):

multiply both side by 15w

OpenStudy (gorv):

\[15w*\frac{ 2 }{ 3w }=15w*(\frac{ 2 }{ 15 }+\frac{ 12 }{ 5w })\]

OpenStudy (gorv):

\[5*2=15w*\frac{ 2 }{ 15 }+15w*\frac{ 12 }{ 5w }\]

OpenStudy (gorv):

\[10=2w+3*12\]

OpenStudy (gorv):

10-36=2w

OpenStudy (gorv):

-26/2=w w=-13